OpenStudy (anonymous):

13. A catapult launches a boulder with an upward velocity of 148 ft/s. The height of the boulder (h) in feet after t seconds is given by the function h = –16t² + 148t + 30. How long does it take the boulder to reach its maximum height? What is the boulder’s maximum height? Round to the nearest hundredth, if necessary. (1 point) A) 9.25 s; 30 ft B) 4.63 s; 640.5 ft C) 4.63 s; 1,056.75 ft D) 4.63 s; 372.25 ft 16. Solve the equation using the Zero Product Property. (2x – 4)(2x – 1) = 0 (1 point) A) 2, –1over2 B) 2, 1over2 C) –2, 2 D) –2, 1over2

OpenStudy (e.cociuba):

1.) 32t + 148 = 0 148 = 32t 4.625 = t t ≈ 4.63 h = -16 • (4.625)^2 + 148 • 4.625 + 30 = 372.25 4.63 s; 372.25 ft ANSWER: D.) 2.) 2x - 4 = 0 and 2x - 1 = 0 x = {4/2, ½} x = {2,½} ANSWER: B.)

OpenStudy (anonymous):

Thanks :)

OpenStudy (e.cociuba):

Np:)

OpenStudy (anonymous):

http://www.jiskha.com/display.cgi?id=1363045983 this explains it pretty well